OpenStudy (anonymous):

Factor completely: 5ab + 3ay + 5b + 3y (5b + 3y)(a + 1) (5b − 3y)(a + 1) (5b + 3y)(a − 1) Prime

jimthompson5910 (jim_thompson5910):

5ab + 3ay + 5b + 3y a(5b + 3y) + 1(5b + 3y) (a+1)(5b + 3y) (5b + 3y)(a+1)

OpenStudy (anonymous):

Simplifying 5ab + 3ay + 5b + 3y = 0 Solving 5ab + 3ay + 5b + 3y = 0 Solving for variable 'a'. Move all terms containing a to the left, all other terms to the right. Add '-5b' to each side of the equation. 5ab + 3ay + 5b + -5b + 3y = 0 + -5b Combine like terms: 5b + -5b = 0 5ab + 3ay + 0 + 3y = 0 + -5b 5ab + 3ay + 3y = 0 + -5b Remove the zero: 5ab + 3ay + 3y = -5b Add '-3y' to each side of the equation. 5ab + 3ay + 3y + -3y = -5b + -3y Combine like terms: 3y + -3y = 0 5ab + 3ay + 0 = -5b + -3y 5ab + 3ay = -5b + -3y Reorder the terms: 5ab + 3ay + 5b + 3y = -5b + 5b + -3y + 3y Combine like terms: -5b + 5b = 0 5ab + 3ay + 5b + 3y = 0 + -3y + 3y 5ab + 3ay + 5b + 3y = -3y + 3y Combine like terms: -3y + 3y = 0 5ab + 3ay + 5b + 3y = 0